Write the equation y - 2 = 4(x + 5) in standard form.
Answer by Cromlix(4381)

I take it you mean in the form
y = mx + c.
y - 2 = 4(x + 5)
y - 2 = 4x + 20
y = 4x + 20 + 2
y = 4x + 22
Hope this helps.
:-)